# Track2Act: Predicting Point Tracks from Internet Videos enables Diverse Zero-shot Robot Manipulation

### v1

@@ Abstract
We seek to learn a generalizable goal-conditioned policy that enables zero-shot robot manipulation — interacting with unseen objects in novel scenes without test-time adaptation. While typical approaches rely on a large amount of demonstration data for such generalization, we propose an approach that leverages web videos to predict plausible interaction plans and learns a task-agnostic transformation to obtain robot actions in the real world. Our framework, Track2Act  predicts tracks of how points in an image should move in future time-steps based on a goal, and can be trained with diverse videos on the web including those of humans and robots manipulating everyday objects. We use these 2D track predictions to infer a sequence of rigid transforms of the object to be manipulated, and obtain robot end-effector poses that can be executed in an open-loop manner. We then refine this open-loop plan by predicting residual actions through a closed loop policy trained with a few embodiment-specific demonstrations. We show that this approach of combining scalably learned track prediction with a residual policy requiring minimal in-domain robot-specific data enables zero-shot robot manipulation, and present a wide array of real-world robot manipulation results across unseen tasks, objects, and scenes. https://homangab.github.io/track2act/111 equal contribution. Correspondence to Homanga Bharadhwaj hbharadh@cs.cmu.edu
Figure 1: Glimpse of some of the diverse robot manipulation capabilities across physical office and kitchen scenes enabled by Track2Act. We learn to predict point tracks from web videos for learning interaction plans that can be used for inferring robot actions in unseen scenarios. This enables a common goal-conditioned policy to perform everyday tasks like closing microwaves, pulling out drawers, flipping open toasters, pouring from jars etc. Columns show first and last images of rollouts from our policy.
@@ 1 Introduction
Robots that can be reliably deployed out-of-the-box in new scenarios have the potential for helping humans in everyday tasks. To realize this vision of generalizable robot manipulation, it is crucial to develop zero-shot execution capabilities i.e. being able to execute a task out-of-the-box without requiring any test-time training through demonstrations or self-practice before solving a specified task. This is an important desiderata for the system to be repeatedly usable without any downtime, and safe to work alongside humans without performing any exploratory actions. We pursue the goal of developing such zero-shot robot manipulation systems that can perform a broad set of everyday tasks. In addition to being deployable zero-shot, to be widely accessible, we aim to make the robot manipulators generalizable to diverse offices, and kitchens in the real world.
@@
Developing zero-shot manipulation capabilities has been attempted by prior works, through behavior cloning on robot interaction datasets [1, 2, 3, 4]. While this approach is in-principle scalable with data, collecting diverse real-world robot interaction data is challenging due to operational constraints. Indeed, recent works that have attempted to scale robot datasets, including cross-robot and cross-domain datasets [5, 6, 1] still suffer from task diversity issues and are mostly restricted to lab-like structured scenarios. Instead of learning a single-policy that can be zero-shot deployed, some recent works aimed at in-the-wild deployment have adopted the method of test-time training [7, 8]. They require either a video of a human performing the task [7] followed by online exploration, or a demo through a robot end-effector held by a human [8]. These approaches are not very convenient for diverse deployments because they require a human to solve the task first, and several hours after that for the robot to learn how to solve that exact task in the exact scene. Thus, such approaches are not zero-shot deployable for new tasks in new scenes.
@@
Our insight to develop an in-the-wild manipulation strategy that is also zero-shot deployable is to factorize a manipulation policy into an interaction-plan that can leverage diverse large-scale video sources on the web of humans and robots manipulating everyday objects and a residual policy that requires a small amount of embodiment-specific robot interaction data. Such a factorized structure is inspired by prior works (e.g. [9]), however, unlike hand-object masks in [9], we instantiate this interaction-plan in an embodiment agnostic-manner by predicting how points in an image of the initial scene move in future frames. This choice of an interaction-plan is more expressive compared to hand-object masks adopted by Bharadhwaj et al. [9], as it directly captures point correspondences across time, while at the same time being easier to compute than full RGB frames [10]. Given an initial image of the scene, a goal image defining the task to be performed, and a random set of points in the initial image, we define the interaction plan to be a 2D trajectory of the locations of the points in future frames, such that the goal is achieved. Importantly, we can train this model purely from the abundantly available human and robot videos on the web without any data specific to the deployment robot embodiment, by using off-the-shelf point-tracking approaches [11] for generating the ground-truth point trajectories. For deployment in a robot’s environment, we can convert the 2D interaction-plan to a sequence of 3D end-effector poses, by having a depth image of the initial scene as an additional input and solving an optimization problem to obtain rigid transforms of the object being manipulated. Finally, with a small amount of embodiment-specific robot interaction data for different tasks ($\sim 400$ trajectories overall), we can learn a goal-conditioned residual policy that corrects for errors in the predicted plan at each time-step and allows for closed-loop deployment.
@@
  • We show how the interaction-plan prediction model can be used for obtaining 3D rigid transforms in a robot’s environment for zero-shot manipulation without using any robot data or online exploration.
  • Given a few ($\sim 400$) embodiment-specific task demonstrations, we show how to learn a goal-conditioned residual policy that can correct for errors in the predicted plan at each time-step. The interaction-plan prediction model combined with the residual policy correction can then be used for zero-shot closed-loop deployment for new tasks in new scenes.
Our real-world robot manipulation results with a Spot robot (highlighted in Fig. 1) show broad generalization across diverse tasks involving unseen objects in unseen scenes, and demonstrate the potential for leveraging easily available passive videos on the web for learning embodiment-agnostic interaction plans. This is significant as it enables zero-shot robot manipulation with a common goal-conditioned policy, that generalizes to unseen tasks without requiring collection of large scale in-domain manipulation datasets.
Figure 2: Illustration of the Track2Act pipeline for learning track prediction from web video datasets, inferring rigid transforms of objects based on the predicted tracks in a robot’s environment, and fine-tuning with a residual policy learned with limited robot data. This approach allows us to learn a single goal-conditioned policy for diverse (unseen) tasks.
@@ 2 Related Works
Understanding interactions from videos. Several computer vision methods have investigated deciphering interactions between hands and objects across various daily activities through curation of large-scale video datasets [12, 13, 14, 15, 16, 17, 18], hand pose estimation [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], object pose estimation [29, 30, 31, 32, 33], interaction hotspot/grasp prediction [34, 35, 36, 37, 38]. Some approaches have investigated generating videos given a description of the task, and often conditioned on a scene [39, 10, 40, 41]. Other works have attempted understanding generic videos by identifying visual correspondences between frames [42, 43]. Recent approaches have made significant advances on this problem by developing general-purpose video tracking systems that can track points specified in a frame, across other frames in the video [44, 11]. Our track prediction model is inspired by these developments, and is based on leveraging the video tracking approaches to generate ground-truth tracks from web videos, and training a model to predict future points tracks of arbitrary points given an initial image and a goal.
@@
Learning Visual Representations for Manipulation.Visual imitation is a promising technique for generalizable robot manipulation [45, 46, 47, 48]. Recent works that have scaled this approach for learning large-scale models for manipulation require extremely high number of expert robot trajectories, often demanding years for collection [2, 3, 1], and still suffer from limited generalization to unseen scenarios for novel objects. Going beyond image observations, prior works have also investigated structured representations like point-clouds [49, 50, 51] and keypoints [52] for manipulation, but are restricted to tasks in structured table-top scenarios. Some of these that predict action in the form of flow-based representations [49, 42] require 3D datasets of robot interactions (often from simulation) which constrain them from generalizable real-world deployments. More recently, Vecerik et al. [53] use point tracking for visual servoing, and the setup requires structured multi-stage definitions of the task and is limited to only minor test-time variations compared to training data. Concurrent work [54] that improves upon [53] by predicting future tracks of points in the current image can learn a policy by combining in-domain human videos with in-domain robot videos. However, the framework is not directly amenable for leveraging web videos because the policy relies on per-step image observations for track prediction. Compared to this, and developed independently, we learn to predict trajectories of arbitrary points from web videos given just an initial image and a goal. We show how we can use these predicted tracks to infer rigid transforms of objects for open-loop execution, and further improve the open-loop plan by predicting residuals over the actions, for closed-loop deployment. This enables much diverse robot manipulation behaviors with a single model, that generalizes to unseen novel objects and scenes in-the-wild.
@@
Leveraging Non-Robot Datasets for Manipulation. One common way of using data beyond robot interactions for efficient learning is to pre-train the visual representations which serve as backbones for the policy models [55, 56, 57, 58, 59] with passive human videos [15, 60] and image data [61]. However, these methods still crucially rely on a lot of in-domain robot data or deployment-time training, and are restricted to learning task-specific policies. Some works that do not require deployment-time training, go beyond visual representations and use curated data of human videos to leverage human hand motion information [62, 63] for learning task-specific policies (instead of a single model across generic tasks). Others that train a single policy across tasks require large in-domain perfectly aligned human-robot data [64, 65, 66] and are not capable of leveraging web data. Towards learning structure more directly related to manipulation from web videos, some works try to predict visual affordances in the form of where to interact in an image, and local information of how to interact [37, 36, 67, 35]. While these could serve as good initializations for a robotic policy, they are not sufficient on their own for accomplishing tasks, and so are typically used in conjunction with online learning, requiring several hours of deployment-time training and robot data [7, 67]. Others learn to predict masks of hand and objects in the scene [9] for conditional behavior cloning and are unable to leverage information of accurate object state changes that is usually ambiguous with a mask. Our work differs from these in terms of predicting an approximate motion of how objects in the scene move in the future through point tracks for the entire trajectory and is zero-shot in terms of not requiring any deployment-time training.
@@ 3 Approach
We aim to develop a zero-shot robot manipulation system, Track2Act  that can scalably leverage diverse video data for generalizable real-world manipulation. Our key insight (Fig. 2) is to have a factorized policy for 1) learning embodiment-agnostic interactions plans of how points in an image of a scene should move in subsequent time-steps to realize a specified goal, followed by 2) inferring robot actions based on the interaction plan through a residual policy. We show how Track2Act  allows us to generalize to diverse scenarios involving unseen tasks and objects, since the prediction model by virtue of being trained on web data generalizes well to new scenes, and the residual policy has a much simpler task of correcting the robot actions derived from the interaction plan.
@@ 3.2 Point Track Prediction from Web Videos
For Track2Act, we instantiate track prediction as a denoising process through a DiT based diffusion model [68]. Let $\mathbf{I_{0}}$ denote the first frame of a video, and $\mathcal{G}$ denote the goal, which we consider to be the last frame of the video. For longer videos, we obtain multiple video clips of 4-5 seconds each for training. Let there be $p$ points in the initial frame to be tracked, such that $P_{0}$ denotes the set of those points and let $H$ be the prediction horizon. $[P_{t}]_{t=1}^{H}$ denotes the future locations of those points in the subsequent time-steps that we want to predict. In the forward diffusion process, all the points $P_{t}$ are corrupted by incrementally adding noise $\mathbf{\epsilon}_{k}$ ($k$ denotes the diffusion time-step), to obtain $\tilde{P}_{t}$, and converging to a unit Gaussian distribution $N(\mathbf{0},\mathbf{I})$. New samples can be generated by reversing the forward diffusion process, by going from Gaussian noise back to the space of point locations. To solve the reverse diffusion process, we need to train a noise predictor $\mathcal{V}_{\theta}(\mathbf{I_{0}},\mathcal{G},P_{0},k)$. We design a DiT Transformer based architecture [68] for $\mathcal{V}_{\theta}$ illustrated visually in Fig. 3. Different from the original DiT model, we condition on embeddings of initial ($z_{0}$) and goal ($z_{g}$) images in addition to that of the diffusion step ($z_{k}$). The input to the Transformer in each batch is a sequence of $p$ tokens corresponding the number of points specified for tracking. The initial $P_{0}$ points are not noisy, as is the convention in training conditional diffusion models on time-series data.
@@
Figure 4: Architecture of the residual policy of Track2Act  that predicts corrections at each time-step over the predicted open-loop plan, and enables closed-loop deployment.
@@ 3.4 Closed-loop Manipulation with Residual Policy Correction
Algorithm 2 Closed-Loop Deployment with Residual Policy Correction
1:procedure Residual Correction($\mathcal{V}_{\theta}(\cdot),\mathbf{I_{0}},\mathcal{G},P_{0},P_{0}^{\text{3D}},\mathbf{K},h,\mathbf{e}_{0}$,$\pi_{\text{res}}$)
2:     $\mathbf{\tau}=\mathcal{V}_{\theta}(\mathbf{I_{0}},\mathcal{G},P_{0})$triangleright Predict future point tracks
3:     $[\mathbf{T}_{t}]_{t=1}^{H}$ = Rigid Transforms($\tau,\mathbf{I_{0}},\mathcal{G},P_{0}^{\text{3D}},\mathbf{K},h$)
4:     for $t\leftarrow 0$ to $H$ do
5:         $\bar{\mathbf{a}}_{t}=\mathbf{T}_{t}\mathbf{e}_{0}$triangleright Infer Open-Loop Plan
6:     for $t\leftarrow 1$ to $H$ do
7:         $\Delta\mathbf{a}_{t:t+h}=\pi_{\text{res}}(\mathbf{I}_{t},\mathcal{G},\mathbf{\tau},[\bar{\mathbf{a}}_{t}]_{t=t}^{t+h})$triangleright Predict residual correction
8:         $\hat{\mathbf{a}}_{t}=\bar{\mathbf{a}}_{t}+\Delta\mathbf{a}_{t}$triangleright Corrected action at time $t$
9:         Execute action $\hat{\mathbf{a}}_{t}$ with the robot $\triangleright$ with IK and gripper action
10:         if t == H then
11:              Gripper Open; Reset end-effector to home

### v2

# Track2Act: Predicting Point Tracks from Internet Videos enables Generalizable Robot Manipulation

@@ Abstract
We seek to learn a generalizable goal-conditioned policy that enables diverse robot manipulation — interacting with unseen objects in novel scenes without test-time adaptation. While typical approaches rely on a large amount of demonstration data for such generalization, we propose an approach that leverages web videos to predict plausible interaction plans and learns a task-agnostic transformation to obtain robot actions in the real world. Our framework, Track2Act  predicts tracks of how points in an image should move in future time-steps based on a goal, and can be trained with diverse videos on the web including those of humans and robots manipulating everyday objects. We use these 2D track predictions to infer a sequence of rigid transforms of the object to be manipulated, and obtain robot end-effector poses that can be executed in an open-loop manner. We then refine this open-loop plan by predicting residual actions through a closed loop policy trained with a few embodiment-specific demonstrations. We show that this approach of combining scalably learned track prediction with a residual policy requiring minimal in-domain robot-specific data enables diverse generalizable robot manipulation, and present a wide array of real-world robot manipulation results across unseen tasks, objects, and scenes. https://homangab.github.io/track2act/ 111 equal contribution. Correspondence to Homanga B. hbharadh@cs.cmu.edu
Keywords:
Diverse Generalizable Manipulation Web Videos
Figure 1: Glimpse of some of the diverse robot manipulation capabilities across physical office and kitchen scenes enabled by our framework. We learn to predict point tracks from web videos for learning interaction plans that can be used for inferring robot actions in unseen scenarios. This enables a common goal-conditioned policy to perform everyday tasks like closing microwaves, pulling out drawers, flipping open toasters, pouring from jars etc. Columns show first and last images of rollouts from our policy.
@@ 1 Introduction
Robots that can be reliably deployed out-of-the-box in new scenarios have the potential for helping humans in everyday tasks. To realize this vision of generalizable robot manipulation, it is crucial to develop direct execution capabilities i.e. being able to execute a task out-of-the-box without requiring any test-time training through demonstrations or self-practice before solving a specified task. This is an important desiderata for the system to be repeatedly usable without any downtime, and safe to work alongside humans without performing any exploratory actions. We pursue the goal of developing such directly executable robot manipulation systems that can perform a broad set of everyday tasks. In addition to being deployable directly, to be widely accessible, we aim to make the robot manipulators generalizable to diverse offices, and kitchens in the real world.
@@
Developing such directly executable manipulation capabilities has been attempted by prior works, through behavior cloning on robot interaction datasets [5, 8, 73, 28]. While this approach is in-principle scalable with data, collecting diverse real-world robot interaction data is challenging due to operational constraints. Indeed, recent works that have attempted to scale robot datasets, including cross-robot and cross-domain datasets [46, 15, 5] still suffer from task diversity issues and are mostly restricted to lab-like structured scenarios. Instead of learning a single-policy that can be directly deployed, some recent works aimed at in-the-wild deployment have adopted the method of test-time training [2, 40]. They require either a video of a human performing the task [2] followed by online exploration, or a demo through a robot end-effector held by a human [40]. These approaches are not very convenient for diverse deployments because they require a human to solve the task first, and several hours after that for the robot to learn how to solve that exact task in the exact scene. Thus, such approaches are not directly deployable for new tasks in new scenes.
@@
Our insight to develop an in-the-wild manipulation strategy that is also directly deployable is to factorize a manipulation policy into an interaction-plan that can leverage diverse large-scale video sources on the web of humans and robots manipulating everyday objects and a residual policy that requires a small amount of embodiment-specific robot interaction data. Such a factorized structure is inspired by prior works (e.g. [4]), however, unlike hand-object masks in [4], we instantiate this interaction-plan in an embodiment agnostic-manner by predicting how points in an image of the initial scene move in future frames. This choice of an interaction-plan is more expressive compared to hand-object masks adopted by Bharadhwaj et al. [4], as it directly captures point correspondences across time, while at the same time being easier to compute than full RGB frames [14]. Given an initial image of the scene, a goal image defining the task to be performed, and a random set of points in the initial image, we define the interaction plan to be a 2D trajectory of the locations of the points in future frames, such that the goal is achieved. Importantly, we can train this model purely from the abundantly available human and robot videos on the web without any data specific to the deployment robot embodiment, by using off-the-shelf point-tracking approaches [29] for generating the ground-truth point trajectories. For deployment in a robot’s environment, we can convert the 2D interaction-plan to a sequence of 3D end-effector poses, by having a depth image of the initial scene as an additional input and solving an optimization problem to obtain rigid transforms of the object being manipulated. Finally, with a small amount of embodiment-specific robot interaction data for different tasks ($\sim 400$ trajectories overall), we can learn a goal-conditioned residual policy that corrects for errors in the predicted plan at each time-step and allows for closed-loop deployment.
@@
  • We show how the interaction-plan prediction model can be used for obtaining 3D rigid transforms in a robot’s environment for direct manipulation without using any robot data or online exploration.
  • Given a few ($\sim 400$) embodiment-specific task demonstrations, we show how to learn a goal-conditioned residual policy that can correct for errors in the predicted plan at each time-step. The interaction-plan prediction model combined with the residual policy correction can then be used for closed-loop deployment for new tasks in new scenes.
Our real-world robot manipulation results with a Spot robot (highlighted in Fig. 1) show broad generalization across diverse tasks involving unseen objects in unseen scenes, and demonstrate the potential for leveraging easily available passive videos on the web for learning embodiment-agnostic interaction plans. This is significant as it enables robot manipulation with a common goal-conditioned policy, that generalizes to unseen tasks without requiring collection of large scale in-domain manipulation datasets.
Figure 2: Illustration of the pipeline for learning track prediction from web video datasets, inferring rigid transforms of objects based on the predicted tracks in a robot’s environment, and fine-tuning with a residual policy learned with limited robot data. This approach allows us to learn a single goal-conditioned policy for diverse (unseen) tasks.
@@ 2 Related Works
Understanding interactions from videos. Several computer vision methods have investigated deciphering interactions between hands and objects across various daily activities through curation of large-scale video datasets [21, 11, 59, 22, 55, 10, 35], hand pose estimation [72, 27, 58, 18, 1, 6, 24, 34, 37, 53], object pose estimation [31, 52, 64, 26, 25], interaction hotspot/grasp prediction [44, 38, 20, 43, 7]. Some approaches have investigated generating videos given a description of the task, and often conditioned on a scene [68, 14, 23, 69]. Other works have attempted understanding generic videos by identifying visual correspondences between frames [19, 36]. Recent approaches have made significant advances on this problem by developing general-purpose video tracking systems that can track points specified in a frame, across other frames in the video [13, 29]. Our track prediction model is inspired by these developments, and is based on leveraging the video tracking approaches to generate ground-truth tracks from web videos, and training a model to predict future points tracks given an initial image and a goal.
@@
Learning Visual Representations for Manipulation.Visual imitation is a promising technique for generalizable robot manipulation [16, 70, 42]. Recent works that have scaled this approach for learning large-scale models for manipulation require extremely high number of expert robot trajectories, often demanding years for collection [8, 73, 5], and still suffer from limited generalization to unseen scenarios for novel objects. Going beyond image observations, prior works have also investigated structured representations like point-clouds [54, 47, 9] and keypoints [51] for manipulation, but are restricted to tasks in structured table-top scenarios. Some of these that predict action in the form of flow-based representations [54, 19] require 3D datasets of robot interactions (often from simulation) which constrain them from generalizable real-world deployments. More recently, Vecerik et al. [60] use point tracking for visual servoing, and the setup requires structured multi-stage definitions of the task and is limited to only minor test-time variations compared to training data. Concurrent work [63] that improves upon [60] by predicting future tracks of points in the current image can learn a policy by combining in-domain human videos with in-domain robot videos. However, the framework is not directly amenable for leveraging web videos because the policy relies on per-step image observations for track prediction. Compared to this, and developed independently, we learn to predict trajectories of arbitrary points from web videos given just an initial image and a goal, we show how we can use these predicted tracks to infer rigid transforms of objects for open-loop execution, and further improve the open-loop plan by predicting residuals over the actions, for closed-loop deployment. This enables much diverse robot manipulation behaviors with a single model, that generalizes to unseen novel objects and scenes in-the-wild.
@@
Leveraging Non-Robot Datasets for Manipulation. One common way of using data beyond robot interactions for efficient learning is to pre-train the visual representations which serve as backbones for the policy models [45, 39, 41, 48, 65] with passive human videos [22, 30] and image data [12]. However, these methods still crucially rely on a lot of in-domain robot data or deployment-time training, and are restricted to learning task-specific policies. Some works that do not require deployment-time training, go beyond visual representations and use curated data of human videos to leverage human hand motion information [50, 56] for learning task-specific policies (instead of a single model across generic tasks). Others that train a single policy across tasks require large in-domain perfectly aligned human-robot data [62, 57, 66] and are not capable of leveraging web data. Towards learning structure more directly related to manipulation from web videos, some works try to predict visual affordances in the form of where to interact in an image, and local information of how to interact [43, 20, 3, 38]. While these could serve as good initializations for a robotic policy, they are not sufficient on their own for accomplishing tasks, and so are typically used in conjunction with online learning, requiring several hours of deployment-time training and robot data [2, 3]. Others learn to predict masks of hand and objects in the scene [4] for conditional behavior cloning and are unable to leverage information of accurate object state changes that is usually ambiguous with a mask. Our work differs from these in terms of predicting an approximate motion of how objects in the scene move in the future through point tracks for the entire trajectory and is directly executable in terms of not requiring any deployment-time training.
@@ 3 Approach
We aim to develop a generalizable robot manipulation system that can scalably leverage diverse video data for generalizable real-world manipulation. Our key insight (Fig. 2) is to have a factorized policy for 1) learning embodiment-agnostic interactions plans of how points in an image of a scene should move in subsequent time-steps to realize a specified goal, followed by 2) inferring robot actions based on the interaction plan through a residual policy. We show how this approach allows us to generalize to diverse scenarios involving unseen tasks and objects, since the prediction model by virtue of being trained on web data generalizes well to new scenes, and the residual policy has a much simpler task of correcting the robot actions derived from the interaction plan.
@@ 3.2 Point Track Prediction from Web Videos
We instantiate track prediction as a denoising process through a DiT based diffusion model [49]. Let $\mathbf{I_{0}}$ denote the first frame of a video, and $\mathcal{G}$ denote the goal, which we consider to be the last frame of the video. For longer videos, we obtain multiple video clips of 4-5 seconds each for training. Let there be $p$ points in the initial frame to be tracked, such that $P_{0}$ denotes the set of those points and let $H$ be the prediction horizon. $[P_{t}]_{t=1}^{H}$ denotes the future locations of those points in the subsequent time-steps that we want to predict. In the forward diffusion process, all the points $P_{t}$ are corrupted by incrementally adding noise $\mathbf{\epsilon}_{k}$ ($k$ denotes the diffusion time-step), to obtain $\tilde{P}_{t}$, and converging to a unit Gaussian distribution $N(\mathbf{0},\mathbf{I})$. New samples can be generated by reversing the forward diffusion process, by going from Gaussian noise back to the space of point locations. To solve the reverse diffusion process, we need to train a noise predictor $\mathcal{V}_{\theta}(\mathbf{I_{0}},\mathcal{G},P_{0},k)$. We design a DiT Transformer based architecture [49] for $\mathcal{V}_{\theta}$ illustrated visually in Fig. 3. Different from the original DiT model, we condition on embeddings of initial ($z_{0}$) and goal ($z_{g}$) images in addition to that of the diffusion step ($z_{k}$). The input to the Transformer in each batch is a sequence of $p$ tokens corresponding the number of points specified for tracking. The initial $P_{0}$ points are not noisy, as is the convention in training conditional diffusion models on time-series data.
@@
Figure 4: Architecture of the residual policy that predicts corrections at each time-step over the predicted open-loop plan, and enables closed-loop deployment.